OpenStudy (anonymous):

A candy company makes chocolates in the shape of solid cones. The original machine for making these chocolates processed chocolate at the rate of 100 kg per hour. The individual chocolate measured 1.5 cm across the base and 2 cm high. Now three new machines are being installed to make three other sizes, listed below. The three new machines will all produce the same number of pieces per hour as the original machine, but they will process chocolate at different rates, because the cones have different volumes. How much chocolate will each new machine process per hour? D = 2.0 cm H=2.0cm

hero (hero):

What are you studying?

hero (hero):

This is calculus, isn't it?

hero (hero):

By the way, I don't see all of the sizes listed.

OpenStudy (anonymous):

I ran out of room. D=2.5 H=3.5 D=3.5 H= 3.5 I attempted this one but I think it is wrong.

OpenStudy (anonymous):

I hope this isn't calculus because I have no idea how to do calculus. The highest I ever got was algebra.

OpenStudy (anonymous):

volume of the chocolate cone made by the first machine V = 1/3 pi r^2 h r = 0.75 V = 1/3 * pi * 0.75^2 * 2 = (1) volume of chocolate cone made by 2nd machine V = 1/3 * pi * 1^2 * 2 = (2) and so on for 3rd and 4th machine.

OpenStudy (anonymous):

then, because the number of chocolate pieces made per hour are the same for all machines 100 kg = volume of 1st machine chocolate piece (1) * T (some fixed constant for pieces of chocolate, and this number doesn't matter because all machines produce the same number of chocolate) 100 kg = volume of 1st machine chocolate piece (1) * T ? kg = volume of 2nd machine chocolate piece (2) * T ? kg = volume of 3rd... (3) * T ? kg = volume of 4th... (4) * T tell me if you have problems understanding this :c